A converging lens of focal length 9.000 сm is 20.0 cm to the left of a diverging lens of focal length -6.00 cm. A coin is plaPart B Express your answer using two significant figures. Vo AED ? di = 22 cm to the left of the converging lens Submit PreviA converging lens of focal length 9.000 cmcm is 20.0 cmcm to the left of a diverging lens of focal length -6.00 cmcm. A coin is placed 19.0 cmcm to the right of the diverging lens.
Answer #1

Solution in the uploaded images

Solution fdem ty = – blon tve ve H 20cm 18cm liverging lens: For for fiz-ben tou= - 19h job- d 1 8.77cm ce u= 20- To the righPpart a From above m, & -8.77 -19 = 0.46 me=w 13.1 - 28.77 -0.455 Total magnification mana m, xm, 0.46x -0.455 elm = -0.209 H